When matt was 11 years old, geom was 15. how many years must pass for matt to be seven years olderthan one-half as old as Geom?
Answer by ankor@dixie-net.com(22740)

When matt was 11 years old, geom was 15.
:
how many years must pass for matt to be seven years older than one-half as old as Geom?
let y = no. of yrs for this to true
11 + y = .5(15+y) + 7
Get rid of the decimal mult both sides by 2
2(11 + y) = 15 + y + 14
22 + 2y = y + 29
2y - y = 29 - 22
y = 7 yrs for this to be true
:
:
check this
11 + 7 = .5(15+7) + 7
18 = .5(22) + 7
18 = 11 + 7